OpenStudy (anonymous):

Solve the quadratic equation x2 - x = 30 by completing the square. A. {-5, 6} B. {-5, -6} C. {4, 5} D. {-4, -5}

OpenStudy (anonymous):

BRB

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Im not sure on how to do these still

OpenStudy (solomonzelman):

x2 is x^2?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

(x-.5)(x-.5)+.25=30 (x-.5)^2=30-.25 (x-.5)^2=29.75 x-.5= √(29.75) x=√(29.75)+.5 x=1/2 √(119) Please ask to explain a step if you don't understand, I want to help you!

OpenStudy (solomonzelman):

karezi-master, you can't have a radical in denominator!

OpenStudy (anonymous):

\[x^{2}-x=30\]

OpenStudy (anonymous):

Thats how the equation is.

OpenStudy (solomonzelman):

you asked to complete the sq... Yes it is! Do u need more explanations? Ask please if u do!

OpenStudy (anonymous):

Don't i do \[(\frac{ -1 }{ 2} )^{2}\]

OpenStudy (anonymous):

Ohh wait i see you use decimals

OpenStudy (anonymous):

\[x ^{2}-x+\frac{ 1 }{ 4 }=30+\frac{ 1 }{4 }\]

OpenStudy (anonymous):

\[\sqrt{x ^{2}}+\sqrt{\frac{ 1 }{ 4 }}=\frac{ 129 }{ 4 }\]

OpenStudy (anonymous):

@SolomonZelman am i doing it right

OpenStudy (mertsj):

30 + 1/4 is 121/4 not 129/4

OpenStudy (solomonzelman):

You know this staff better than me! Are you a math teacher? Where do you teach?

OpenStudy (anonymous):

ohh yeah messed up on that

OpenStudy (mertsj):

\[x^2-x=30\] \[x^2-x+\frac{1}{4}=30+\frac{1}{4}\] \[(x-\frac{1}{2})^2=\frac{121}{4}\] \[x-\frac{1}{2}=\pm \frac{11}{2}\] \[x=\frac{11}{2}+\frac{1}{2}=\frac{12}{2}=6\] \[x=\frac{-11}{2}+\frac{1}{2}=\frac{-10}{2}=-5\]

OpenStudy (solomonzelman):

I was doing it right, but MISCALCULATED in my head! My bad....

OpenStudy (anonymous):

\[x ^{2}+12+20=0\]

OpenStudy (anonymous):

\[x ^{2}+12x=20\]

OpenStudy (anonymous):

imma try it my self see if im right

OpenStudy (anonymous):

\[(\frac{ 12 }{ 2})^{2}=36\]

OpenStudy (anonymous):

\[x ^{2}+12x+36=-20+36\]

OpenStudy (anonymous):

\[\sqrt{ x ^{2}}+\sqrt{36}=\sqrt{16}\]

OpenStudy (anonymous):

\[x +6=4\]

OpenStudy (anonymous):

i get (2,10) @Mertsj

OpenStudy (mertsj):

How did the 20 all of a sudden turn into -20?

OpenStudy (anonymous):

I forgot to switch signs when i subtracted 20 to both sides

OpenStudy (mertsj):

There are 2 square roots of 16

OpenStudy (anonymous):

yeah +4 and -4

OpenStudy (anonymous):

soo the other equation is 4+6 and -4+6

OpenStudy (mertsj):

\[x^2+12x=-20\] \[x^2+12x+36=-20+36\] \[(x+6)^2=16\] \[x+6=\pm4\] \[x=4-6 or -4-6\] \[x=-2 or x = -10\]

OpenStudy (anonymous):

Why does the 6 go to negative?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

x+6=4 Can you solve that equation?

OpenStudy (anonymous):

Ohhhh hahaha XD ok Thanks man

OpenStudy (mertsj):

yw